# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 3.4 Self-Improvement
As LLM agents evolve, recent research increasingly emphasizes RL as a mechanism for ongoing reflection, enabling agents to learn from their own mistakes across planning, reasoning, tool use, and memory [129]. Rather than relying exclusively on data-driven training phases or static reward models, these systems incorporate iterative, self-generated feedback loops, ranging from prompt-level heuristics to fully fledged RL controllers, to guide agents toward continual self-improvement.
@@
Initial methods in this vein leveraged prompt-based heuristics, sometimes referred to as verbal reinforcement learning, where agents generate an answer, linguistically reflect on its potential errors, and subsequently produce a refined solution, all within a single inferential pass without gradient updates. Prominent examples include Reflexion [130], Self-refine [131], CRITIC [132], and Chain-of-Verification [133]. For instance, the Self-Refine [131] protocol directs an LLM to iteratively polish its output using three distinct prompts for generation, feedback, and refinement, proving effective across domains like reasoning and programming. To enhance the efficacy and robustness of such self-reflection, several distinct strategies have been developed: (I) multiple sampling, which involves generating multiple output rollouts by sampling from the model’s distribution. By aggregating critiques or solutions from multiple attempts, the agent can improve the consistency and quality of its self-reflection. This method has been widely studied in works like If-or-Else [134], UALA [135] and Multi-agent Verification [136]. This approach is conceptually analogous to test-time scaling techniques, so we refer the reader to [109] for more details; (II) structured reflection workflows, rather than prompting for a monolithic reflection on a final answer, prescribe a more dedicated and granular workflow. For example, Chain-of-Verification [133] manually decomposes the process into distinct “Retrieving, Rethinking, and Revising” stages; (III) external guidance, which grounds the reflection process in verifiable, objective feedback by incorporating external tools. These tools include code interpreter as seen in Self-Debugging [137], CAD modeling programs in Luban [138], mathematical calculators in T1 [139], step-wise reward models [140], and others [132].
@@
While verbal self-correction offers a potent inference-time technique, its improvements are ephemeral and confined to a single session. To instill a more durable and generalized capability for self-improvement, subsequent research has employed RL with gradient-based updates to internalize these reflective feedback loops directly into the model’s parameters and to fundamentally enhance the model’s inherent ability to identify and correct its own errors. This paradigm has been applied across multiple domains. For instance, KnowSelf [141] leverages DPO and RPO [142] to enhance agents’ self-reflection capabilities in text-based game environments, while Reflection-DPO [143] focuses on user–agent interaction scenarios, enabling agents to better infer user intent through reflective reasoning. DuPo [144] employs RL with dual-task feedback to enable annotation-free optimization, enhancing LLM agents’ self-correction across translation, reasoning, and reranking tasks. SWEET-RL [145] and ACC-Collab [146] adopt a slightly different setting from the above works: they train an external critic model to provide higher-quality revision suggestions for the actor agent’s actions. Nonetheless, the underlying principle remains closely aligned.
@@
Moving toward full agentic autonomy, the third and most advanced class of models combines reflection, reasoning, and task generation into a self-sustaining loop, enabling unbounded self-improvement without human-labeled data. These methods can be distinguished by the architecture of their learning loops: (I) Self-play and search-guided refinement, which emulates classic RL paradigms like AlphaZero. R-Zero [147], for instance, employs a Monte Carlo Tree Search (MCTS) to explore a reasoning tree, using the search results to iteratively train both a policy LLM (the actor) and a value LLM (the critic) entirely from scratch. Similarly, the ISC framework [148] operationalizes a cycle of "Imagination, Searching, and Criticizing," where the agent generates potential solution paths, uses a search algorithm to explore them, and applies a critic to refine its reasoning strategy before producing a final answer. (II) Execution-guided curriculum generation, where the agent creates its own problems and learns from verifiable outcomes. Absolute Zero [149] exemplifies this by proposing its own tasks, attempting solutions, verifying them via execution, and using the outcome-based reward to refine its policy. Similarly, Self-Evolving Curriculum [150] enhances this process by framing problem selection itself as a non-stationary bandit task, allowing the agent to strategically generate a curriculum that maximizes its learning gains over time. TTRL [151] applies this principle for on-the-fly adaptation to a single problem. At test time, it uses execution-based rewards to rapidly fine-tune a temporary copy of the agent’s policy for the specific task at hand; this specialized policy is then used to generate the final answer before being discarded. Though differing in whether the learning is permanent or ephemeral, all these methods underscore a powerful, unified strategy: harnessing execution-based feedback to autonomously guide the agent’s reasoning process. ALAS [152] constructs an autonomous pipeline that crawls web data, distills it into training signals, and continuously fine-tunes LLMs, thereby enabling self-training and self-evolution without manual dataset curation. (III) Collective bootstrapping, where learning is accelerated by aggregating shared experience. SiriuS [153], for example, constructs and augments a live repository of successful reasoning trajectories from multi-agent interactions, using this growing knowledge base to bootstrap its own training curriculum. MALT [154] shares a similar motivation, yet its implementation is limited to a three-agent setup. Nevertheless, all these methods define feedback loops that are internally generated and continuously evolving, representing a significant step toward truly autonomous agents.
@@ 3.5 Reasoning
Reasoning in large language models can be broadly categorized into fast reasoning and slow reasoning, following the dual-process cognitive theory [155]. Fast reasoning corresponds to rapid, heuristic-driven inference with minimal intermediate steps, while slow reasoning emphasizes deliberate, structured, and multi-step reasoning. Understanding the trade-offs between these two paradigms is crucial for designing models that balance efficiency and accuracy in complex problem-solving.
@@
Fast reasoning models operate in a manner analogous to System 1 [2] cognition: quick, intuitive, and pattern-driven. They generate immediate responses without explicit step-by-step deliberation, excelling in tasks that prioritize fluency, efficiency, and low latency. Most conventional LLMs fall under this category, where reasoning is implicitly encoded in next-token prediction [13, 156]. However, this efficiency comes at the cost of systematic reasoning, making these models more vulnerable to factual errors, biases, and shallow generalization.
To address the severe hallucination problems in fast reasoning, current research has largely focused on various direct approaches. Some studies attempt to mitigate errors and hallucinations in the next-token prediction paradigm by leveraging internal mechanisms [157, 158, 159] or by simulating human-like cognitive reasoning. Other works propose introducing both external and internal confidence estimation methods [160, 161] to identify more reliable reasoning paths. However, constructing such external reasoning frameworks often risks algorithmic adaptivity issues and can easily fall into the complexity trap.
@@
In contrast, slow reasoning models are designed to emulate System 2 cognition [2] by explicitly producing intermediate reasoning traces. Techniques such as chain-of-thought prompting, multi-step verification [162], and reasoning-augmented reinforcement learning allow these models to engage in deeper reflection and achieve greater logical consistency. While slower in inference due to extended reasoning trajectories, they achieve higher accuracy and robustness in knowledge-intensive tasks such as mathematics, scientific reasoning, and multi-hop question answering [163]. Representative examples include OpenAI’s o1 [30] and o3 series [32], DeepSeek-R1 [31], as well as methods that incorporate dynamic test-time scaling [164, 165, 166, 167] or reinforcement learning [168, 169, 170, 171, 172, 173] for reasoning.
Modern slow reasoning exhibits output structures that differ substantially from fast reasoning. These include a clear exploration and planning structure, frequent verification and checking behaviors, and generally longer inference lengths and times. Past work has explored diverse patterns for constructing long-chain reasoning outputs. Some methods—Macro-o1, HuatuoGPT-o1, and AlphaZero—have attempted to synthesize long chains-of-thought via structured, agentic search [174, 175, 176]. Other approaches focus on generating long-CoT datasets that embody specific deliberative or reflective thinking patterns; examples include HiICL-MCTS, LLaVA-CoT, rStar-Math, and ReasonFlux [177, 178, 179, 180]. With the progress of pretrained foundation models, more recent work has shifted toward self-improvement paradigms—frequently instantiated with reinforcement learning—to further enhance models’ reasoning capabilities [168, 169].
@@
The dichotomy between fast and slow reasoning highlights an open challenge in agentic reasoning: how to employ reinforcement learning for reliably training slow-thinking reasoning capabilities in agentic scenarios. Reinforcement learning in agentic scenarios faces greater challenges in training stability, such as ensuring compatibility with diverse environments. Agentic reasoning itself is also susceptible to overthinking problems. Purely fast models may overlook critical reasoning steps, while slow models often suffer from excessive latency or overthinking behaviors, such as unnecessarily long chains of thought. Emerging approaches seek hybrid strategies [181] that combine the efficiency of fast reasoning with the rigor of slow reasoning [182, 183, 184, 185]. For instance, adaptive test-time scaling allows a model to decide whether to respond quickly or to engage in extended deliberation depending on task complexity. Developing such cognitively-aligned mechanisms is a key step toward building reasoning agents that are both efficient and reliable.
@@ 3.6 Perception
By bridging visual perception with linguistic abstraction, Large Vision–Language Models (LVLMs) have demonstrated unprecedented capabilities for perceiving and understanding multimodal content [186, 187, 188, 189, 190, 191, 192, 193]. Central to this progress is the incorporation of explicit reasoning mechanisms into multimodal learning frameworks [194, 195], moving beyond passive perception toward active visual cognition [196]. RL has emerged as a powerful paradigm for this purpose, enabling the alignment of vision–language–action models with complex, multi-step reasoning objectives that go beyond the constraints of supervised next-token prediction [197, 198].
@@
Multimodal content often requires nuanced, context-dependent interpretation. Inspired by the remarkable success of RL in enhancing reasoning within LLMs [31, 199], researchers have increasingly sought to transfer these gains to multimodal learning [200, 201]. Early efforts focused on preference-based RL to strengthen the Chain-of-Thought (CoT) reasoning ability of MLLMs [202, 203, 204]. Visual-RFT [205] and Reason-RFT [206] directly apply GRPO to the vision domain, adaptively incorporating vision-specific metrics such as IoU as verifiable reward signals, while STAR-R1 [207] extended this idea by introducing partial rewards tailored for visual GRPO. Building upon this, a series of approaches—Vision-R1 [208], VLM-R1 [200], LMM-R1 [201], and MM-Eureka [209]—developed specialized policy optimization algorithms designed to incentivize step-wise visual reasoning, demonstrating strong performance even on smaller 3B-parameter models. SVQA-R1 [210] introduced Spatial-GRPO, a novel groupwise RL method that enforces view-consistent and transformation-invariant objectives. Visionary-R1 [211] enforces image captioning as a prerequisite step before reasoning, mitigating shortcut exploitation during reinforcement finetuning. A line of curriculum-learning methods have also been proposed to ease and smooth the RL training process of vision reinforcement finetuning [212, 213, 214, 215, 203]. R1-V [213] introduces VLM-Gym and trains G0/G1 models via scalable, pure RL self-evolution with a perception-enhanced cold start, yielding emergent perception–reasoning synergy across diverse visual tasks. R1-Zero [216] shows that even simple rule-based rewards can induce self-reflection and extended reasoning in non-SFT models, surpassing supervised baselines. PAPO [63] proposes a perception-aware policy optimization framework that augments RLVR methods with an implicit perception KL loss and double-entropy regularization, while [217] proposes a summarize-and-then-reason framework under RL training to mitigate visual hallucinations and improve reasoning without dense human annotations. Collectively, these approaches demonstrate that R1-style RL can be successfully transferred to the vision domain, provided that well-designed, verifiable reward metrics are used—yielding significant improvements in performance, robustness, and out-of-distribution generalization.
@@
More recent work explores another key advantage of RL: moving beyond the formulation of tasks as passive perception, where static, verifiable rewards are computed only on the text-based outputs of LVLMs. Instead, RL can be used to incentivize active cognition over multimodal content—treating visual representations as manipulable and verifiable intermediate thoughts. This paradigm empowers models not merely to “look and answer,” but to actively see, manipulate, and reason with visual information as part of a multi-step cognitive process [196].
@@
To advance from passive perception to active visual cognition, a key direction is enabling LVLMs to repeatedly look back and query the image while generating their reasoning process. This is achieved through grounding [218, 219], which anchors each step of the generated chain-of-thought (CoT) to specific regions of the multimodal input—facilitating more valid and verifiable reasoning by explicitly linking text with corresponding visual regions.
To begin with, GRIT [220] interleaves bounding-box tokens with textual CoT and uses GRPO with both verifiable rewards and bounding-box correctness as supervision. [221] introduces a simple point-and-copy mechanism that allows the model to dynamically retrieve relevant image regions throughout the reasoning process. Ground-R1 [222] and BRPO [223] highlight evidence regions (via IoU-based or reflection rewards) prior to text-only reasoning, while DeepEyes [88] demonstrates that end-to-end RL can naturally induce such grounding behaviors. Chain-of-Focus further refines this approach by grounding CoT steps followed by zooming in operations.
@@
Another promising direction for enabling active perception is to frame visual cognition as an agentic process, where external tools, code snippets, and runtime environments assist the model’s cognitive workflow [224, 225]. For instance, VisTA [226] and VTool-R1 [227] focus on teaching models how to select and use visual tools through RL, while OpenThinkIMG [228] provides standardized infrastructure for training models to “think with images.” Finally, Visual-ARFT [205] leverages RL to facilitate tool creation, harnessing the code-generation capabilities of MLLMs to dynamically extend their perceptual toolkit. Pixel Reasoner [89] expands the model’s action space with operations such as crop, erase, and paint, and introduces curiosity-driven rewards to discourage premature termination of exploration.
@@
In addition to grounding and tool use, humans employ one of their most powerful cognitive abilities—imagination—to produce sketches or diagrams that aid problem-solving. Inspired by this, researchers have begun equipping LVLMs with the ability to generate sketches or images interleaved with chain-of-thought (CoT) reasoning, enabling models to externalize intermediate representations and reason more effectively [229, 230, 231]. Visual Planning [229] proposes to use imagined image rollouts only as the CoT images thinking, using downstream task success as the reward signal. GoT-R1 [232] applies RL within the Generation-CoT framework, allowing models to autonomously discover semantic–spatial reasoning plans before producing the image. Similarly, T2I-R1 [233] explicitly decouples the process into a semantic-level CoT for high-level planning and a token-level CoT for patch-wise pixel generation, jointly optimizing both stages with RL.
@@
RL has also been extended beyond vision–language models to a diverse range of modalities, including audio. Within the audio–language domain, we categorize RL applications into two broad classes. (1) Reasoning enhancement for large audio–language models: RL is leveraged to guide models in producing structured, step-by-step reasoning chains for tasks such as audio question answering and logical inference [234, 235, 236, 236, 234]. (2) Fine-grained component optimization in speech synthesis (TTS): RL is employed to directly refine system components—for example, improving duration predictors—using perceptual quality metrics such as speaker similarity and word error rate as reward signals, thereby yielding more natural and intelligible speech [237]. Some other works such as EchoInk-R1 [238] further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization.

### v4

@@ 3.4 Self-Improvement
As LLM agents evolve, recent research increasingly emphasizes RL as a mechanism for ongoing reflection, enabling agents to learn from their own mistakes across planning, reasoning, tool use, and memory (ang Gao et al., 2025). Rather than relying exclusively on data-driven training phases or static reward models, these systems incorporate iterative, self-generated feedback loops, ranging from prompt-level heuristics to fully fledged RL controllers, to guide agents toward continual self-improvement.
@@
Initial methods in this vein leveraged prompt-based heuristics, sometimes referred to as verbal reinforcement learning, where agents generate an answer, linguistically reflect on its potential errors, and subsequently produce a refined solution, all within a single inferential pass without gradient updates. Prominent examples include Reflexion (Shinn et al., 2023), Self-refine (Madaan et al., 2023), CRITIC (Gou et al., 2024), and Chain-of-Verification (He et al., 2024). For instance, the Self-Refine (Madaan et al., 2023) protocol directs an LLM to iteratively polish its output using three distinct prompts for generation, feedback, and refinement, proving effective across domains like reasoning and programming. To enhance the efficacy and robustness of such self-reflection, several distinct strategies have been developed: (I) multiple sampling, which involves generating multiple output rollouts by sampling from the model’s distribution. By aggregating critiques or solutions from multiple attempts, the agent can improve the consistency and quality of its self-reflection. This method has been widely studied in works like If-or-Else (Li et al., 2024b), UALA (Han et al., 2024) and Multi-agent Verification (Lifshitz et al., 2025). This approach is conceptually analogous to test-time scaling techniques, so we refer the reader to (Pignatelli et al., 2024) for more details; (II) structured reflection workflows, rather than prompting for a monolithic reflection on a final answer, prescribe a more dedicated and granular workflow. For example, Chain-of-Verification (He et al., 2024) manually decomposes the process into distinct “Retrieving, Rethinking, and Revising” stages; (III) external guidance, which grounds the reflection process in verifiable, objective feedback by incorporating external tools. These tools include code interpreters, as seen in Self-Debugging (Chen et al., 2024f), CAD modeling programs in Luban (Guo et al., 2024), mathematical calculators in T1 (Kang et al., 2025b), step-wise reward models (Xiong et al., 2025), and tool-interactive critiquing mechanisms (Gou et al., 2024).
@@
While verbal self-correction offers a potent inference-time technique, its improvements are ephemeral and confined to a single session. To instill a more durable and generalized capability for self-improvement, subsequent research has employed RL with gradient-based updates to internalize these reflective feedback loops directly into the model’s parameters and to fundamentally enhance the model’s inherent ability to identify and correct its own errors. This paradigm has been applied across multiple domains. For instance, KnowSelf (Qiao et al., 2025) leverages DPO and RPO (Pang et al., 2024) to enhance agents’ self-reflection capabilities in text-based game environments, while Reflection-DPO (Patel et al., 2025) focuses on user–agent interaction scenarios, enabling agents to better infer user intent through reflective reasoning. DuPo (She et al., 2025) employs RL with dual-task feedback to enable annotation-free optimization, enhancing LLM agents’ self-correction across translation, reasoning, and reranking tasks. SWEET-RL (Zhou et al., 2025e) and ACC-Collab (Estornell et al., 2025b) adopt a slightly different setting from the above works: they train an external critic model to provide higher-quality revision suggestions for the actor agent’s actions. Nonetheless, the underlying principle remains closely aligned.
@@
Moving toward full agentic autonomy, the third and most advanced class of models combines reflection, reasoning, and task generation into a self-sustaining loop, enabling unbounded self-improvement without human-labeled data. These methods can be distinguished by the architecture of their learning loops: (I) Self-play and search-guided refinement, which emulates classic RL paradigms like AlphaZero. R-Zero (Huang et al., 2025a), for instance, employs a Monte Carlo Tree Search (MCTS) to explore a reasoning tree, using the search results to iteratively train both a policy LLM (the actor) and a value LLM (the critic) entirely from scratch. Similarly, the ISC framework (Tian et al., 2024) operationalizes a cycle of "Imagination, Searching, and Criticizing," where the agent generates potential solution paths, uses a search algorithm to explore them, and applies a critic to refine its reasoning strategy before producing a final answer. (II) Execution-guided curriculum generation, where the agent creates its own problems and learns from verifiable outcomes. Absolute Zero (Zhao et al., 2025a) exemplifies this by proposing its own tasks, attempting solutions, verifying them via execution, and using the outcome-based reward to refine its policy. Similarly, Self-Evolving Curriculum (Chen et al., 2025f) enhances this process by framing problem selection itself as a non-stationary bandit task, allowing the agent to strategically generate a curriculum that maximizes its learning gains over time. TTRL (Zuo et al., 2025) applies this principle for on-the-fly adaptation to a single problem. At test time, it uses execution-based rewards to rapidly fine-tune a temporary copy of the agent’s policy for the specific task at hand; this specialized policy is then used to generate the final answer before being discarded. Though differing in whether the learning is permanent or ephemeral, all these methods underscore a powerful, unified strategy: harnessing execution-based feedback to autonomously guide the agent’s reasoning process. ALAS (Atreja, 2025) constructs an autonomous pipeline that crawls web data, distills it into training signals, and continuously fine-tunes LLMs, thereby enabling self-training and self-evolution without manual dataset curation. (III) Collective bootstrapping, where learning is accelerated by aggregating shared experience. SiriuS (Zhao et al., 2025e), for example, constructs and augments a live repository of successful reasoning trajectories from multi‑agent interactions, using this growing knowledge base to bootstrap its own training curriculum. MALT (Motwani et al., 2025) shares a similar motivation, yet its implementation is limited to a three-agent setup. Nevertheless, all these methods define feedback loops that are internally generated and continuously evolving, representing a significant step toward truly autonomous agents.
@@ 3.5 Reasoning
Reasoning in large language models can be broadly categorized into fast reasoning and slow reasoning, building on the dual-process cognitive theory (Kahneman, 2011; Kahneman & Tversky, 1974; Stanovich & West, 2000), as discussed in recent surveys (Ke et al., 2025; Kumar et al., 2025). Fast reasoning corresponds to rapid, heuristic-driven inference with minimal intermediate steps, while slow reasoning emphasizes deliberate, structured, and multi-step reasoning. Understanding the trade-offs between these two paradigms is crucial for designing models that balance efficiency and accuracy in complex problem-solving.
@@
Fast reasoning models operate in a manner analogous to System 1 (Li et al., 2025r) cognition: quick, intuitive, and pattern-driven. They generate immediate responses without explicit step-by-step deliberation, excelling in tasks that prioritize fluency, efficiency, and low latency. Most conventional LLMs fall under this category, where reasoning is implicitly encoded in next-token prediction (Shao et al., 2024b; Yang et al., 2024a). However, this efficiency comes at the cost of systematic reasoning, making these models more vulnerable to factual errors, biases, and shallow generalization.
To address the severe hallucination problems in fast reasoning, current research has largely focused on various direct approaches. Some studies attempt to mitigate errors and hallucinations in the next-token prediction paradigm by leveraging internal mechanisms (Wang et al., 2023b; Yao et al., 2023a; Besta et al., 2024) or by simulating human-like cognitive reasoning. Other works propose introducing both external and internal confidence estimation methods (Lightman et al., 2023; Wang et al., 2024d) to identify more reliable reasoning paths. However, constructing such external reasoning frameworks often risks algorithmic adaptivity issues and can easily fall into the complexity trap.
@@
In contrast, slow reasoning models are designed to emulate System 2 cognition (Li et al., 2025r) by explicitly producing intermediate reasoning traces. Techniques such as chain-of-thought prompting, multi-step verification (Qin et al., 2024a), and reasoning-augmented reinforcement learning allow these models to engage in deeper reflection and achieve greater logical consistency. While slower in inference due to extended reasoning trajectories, they achieve higher accuracy and robustness in knowledge-intensive tasks such as mathematics, scientific reasoning, and multi-hop question answering (Chu et al., 2025a). Representative examples include OpenAI’s o1 (OpenAI et al., 2024) and o3 series (OpenAI Team, 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), as well as methods that incorporate dynamic test-time scaling (Aggarwal & Welleck, 2025; Zhang et al., 2024a; Xu et al., 2025a; Yao et al., 2023a) or reinforcement learning (Zeng et al., 2025c; Yu et al., 2025e; Wang et al., 2025k; Liang et al., 2025a; b; Yue et al., 2025a) for reasoning.
Modern slow reasoning exhibits output structures that differ substantially from fast reasoning. These include a clear exploration and planning structure, frequent verification and checking behaviors, and generally longer inference lengths and times. Past work has explored diverse patterns for constructing long-chain reasoning outputs. Some methods—Macro-o1, HuatuoGPT-o1, and AlphaZero—have attempted to synthesize long chains-of-thought via structured, agentic search (Zhao et al., 2024; Chen et al., 2024c; b). Other approaches focus on generating long-CoT datasets that embody specific deliberative or reflective thinking patterns; examples include HiICL-MCTS, LLaVA-CoT, rStar-Math, and ReasonFlux (Wu et al., 2025d; Xu et al., 2025b; Guan et al., 2025a; Yang et al., 2025d). Recent approaches that perform reasoning in the latent space leverage latent representations to conduct parallel reasoning and explore diverse reasoning trajectories (Zhang et al., 2025x; Hao et al., 2024). With the progress of pretrained foundation models, more recent work has shifted toward self-improvement paradigms—frequently instantiated with reinforcement learning—to further enhance models’ reasoning capabilities (Zeng et al., 2025c; Yu et al., 2025e).
@@
The dichotomy between fast and slow reasoning highlights an open challenge in agentic reasoning: how to employ reinforcement learning for reliably training slow-thinking reasoning capabilities in agentic scenarios. Reinforcement learning in agentic scenarios faces greater challenges in training stability, such as ensuring compatibility with diverse environments. Agentic reasoning itself is also susceptible to overthinking. Purely fast models may overlook critical reasoning steps, while slow models often suffer from excessive latency or overthinking behaviors, such as unnecessarily long chains of thought. Emerging approaches seek hybrid strategies (Yang et al., 2025a) that combine the efficiency of fast reasoning with the rigor of slow reasoning (Yang et al., 2025g; Hou et al., 2025; Li et al., 2025q; Chen et al., 2025g). For instance, adaptive test-time scaling allows a model to decide whether to respond quickly or to engage in extended deliberation depending on task complexity. Developing such cognitively aligned mechanisms is a key step toward building reasoning agents that are both efficient and reliable.
Figure 5: A conceptual overview of how RL empowers agentic LLMs across six core capabilities. The central panel summarizes the capability taxonomy, while the side panels illustrate representative RL mechanisms and interaction patterns. Listed methods are illustrative rather than exhaustive; see the main text for details.
@@ 3.6 Perception
By bridging visual perception with linguistic abstraction, Large Vision–Language Models (LVLMs) have demonstrated unprecedented capabilities for perceiving and understanding multimodal content  (Team et al., 2023; Liu et al., 2023a; Wang et al., 2024e; Li et al., 2024d; Chen et al., 2024j; OpenAI, 2023; Zhang et al., 2025p; 2024b). Central to this progress is the incorporation of explicit reasoning mechanisms into multimodal learning frameworks (Shao et al., 2024a; Zhang et al., 2023), moving beyond passive perception toward active visual cognition (Su et al., 2025c). RL has emerged as a powerful paradigm for this purpose, enabling the alignment of vision–language–action models with complex, multi-step reasoning objectives that go beyond the constraints of supervised next-token prediction (Zhou et al., 2025a; Wu et al., 2025h).
@@
Multimodal content often requires nuanced, context-dependent interpretation. Inspired by the remarkable success of RL in enhancing reasoning within LLMs  (DeepSeek-AI et al., 2025; Team et al., 2025b), researchers have increasingly sought to transfer these gains to multimodal learning  (Shen et al., 2025a; Peng et al., 2025). Early efforts focused on preference-based RL to strengthen the Chain-of-Thought (CoT) reasoning ability of MLLMs (Wang et al., 2024g; Dong et al., 2025d; Zhu et al., 2025b). Visual-RFT (Liu et al., 2025v) and Reason-RFT (Tan et al., 2025a) directly apply GRPO to the vision domain, adaptively incorporating vision-specific metrics such as IoU as verifiable reward signals, while STAR-R1 (Li et al., 2025t) extended this idea by introducing partial rewards tailored for visual GRPO. Building upon this, a series of approaches—Vision-R1 (Huang et al., 2025c), VLM-R1 (Shen et al., 2025a), LMM-R1 (Peng et al., 2025), and MM-Eureka (Meng et al., 2025)—developed specialized policy optimization algorithms designed to incentivize step-wise visual reasoning, demonstrating strong performance even on smaller 3B-parameter models. SVQA-R1 (Wang & Ling, 2025) introduced Spatial-GRPO, a novel groupwise RL method that enforces view-consistent and transformation-invariant objectives. Visionary-R1 (Xia et al., 2025a) enforces image captioning as a prerequisite step before reasoning, mitigating shortcut exploitation during reinforcement finetuning. A line of curriculum-learning methods have also been proposed to ease and smooth the RL training process of vision reinforcement finetuning (Yang et al., 2025c; Chen et al., 2025b; Zhan et al., 2025; Guo et al., 2025d; Dong et al., 2025d). R1-V (Chen et al., 2025b) introduces VLM-Gym and trains G0/G1 models via scalable, pure RL self-evolution with a perception-enhanced cold start, yielding emergent perception–reasoning synergy across diverse visual tasks. R1-Zero (Zhou et al., 2025d) shows that even simple rule-based rewards can induce self-reflection and extended reasoning in non-SFT models, surpassing supervised baselines. PAPO (Wang et al., 2025u) proposes a perception-aware policy optimization framework that augments RLVR methods with an implicit perception KL loss and double-entropy regularization, while Li et al. (2025s) proposes a summarize-and-then-reason framework under RL training to mitigate visual hallucinations and improve reasoning without dense human annotations. Collectively, these approaches demonstrate that R1-style RL can be successfully transferred to the vision domain, provided that well-designed, verifiable reward metrics are used—yielding significant improvements in performance, robustness, and out-of-distribution generalization.
@@
More recent work explores another key advantage of RL: moving beyond the formulation of tasks as passive perception, where static, verifiable rewards are computed only on the text-based outputs of LVLMs. Instead, RL can be used to incentivize active cognition over multimodal content—treating visual representations as manipulable and verifiable intermediate thoughts. This paradigm empowers models not merely to “look and answer,” but to actively see, manipulate, and reason with visual information as part of a multi-step cognitive process (Su et al., 2025c).
@@
To advance from passive perception to active visual cognition, a key direction is enabling LVLMs to repeatedly look back and query the image while generating their reasoning process. This is achieved through grounding (Nagaraja et al., 2016; Mao et al., 2016), which anchors each step of the generated chain-of-thought (CoT) to specific regions of the multimodal input—facilitating more valid and verifiable reasoning by explicitly linking text with corresponding visual regions.
To begin with, GRIT (Fan et al., 2025c) interleaves bounding-box tokens with textual CoT and uses GRPO with both verifiable rewards and bounding-box correctness as supervision.  Chung et al. (2025) introduces a simple point-and-copy mechanism that allows the model to dynamically retrieve relevant image regions throughout the reasoning process. Ground-R1 (Cao et al., 2025a) and BRPO (Chu et al., 2025c) highlight evidence regions (via IoU-based or reflection rewards) prior to text-only reasoning, while DeepEyes (Zheng et al., 2025g) demonstrates that end-to-end RL can naturally induce such grounding behaviors. Chain-of-Focus further refines this approach by grounding CoT steps followed by zooming in operations.
@@
Another promising direction for enabling active perception is to frame visual cognition as an agentic process, where external tools, code snippets, and runtime environments assist the model’s cognitive workflow (Gupta & Kembhavi, 2023; Zhao et al., 2025d). For instance, VisTA (Huang et al., 2025d) and VTool-R1 (Wu et al., 2025g) focus on teaching models how to select and use visual tools through RL, while OpenThinkIMG (Su et al., 2025b) provides standardized infrastructure for training models to “think with images.” Finally, Visual-ARFT (Liu et al., 2025v) leverages RL to facilitate tool creation, harnessing the code-generation capabilities of MLLMs to dynamically extend their perceptual toolkit. Pixel Reasoner (Su et al., 2025a) expands the model’s action space with operations such as crop, erase, and paint, and introduces curiosity-driven rewards to discourage premature termination of exploration.
@@
In addition to grounding and tool use, humans employ one of their most powerful cognitive abilities—imagination—to produce sketches or diagrams that aid problem-solving. Inspired by this, researchers have begun equipping LVLMs with the ability to generate sketches or images interleaved with chain-of-thought (CoT) reasoning, enabling models to externalize intermediate representations and reason more effectively (Xu et al., 2025e; Fang et al., 2025a; Li et al., 2025c). Visual Planning (Xu et al., 2025e) proposes to use imagined image rollouts only as the CoT images thinking, using downstream task success as the reward signal. GoT-R1 (Duan et al., 2025) applies RL within the Generation-CoT framework, allowing models to autonomously discover semantic–spatial reasoning plans before producing the image. Similarly, T2I-R1 (Jiang et al., 2025b) explicitly decouples the process into a semantic-level CoT for high-level planning and a token-level CoT for patch-wise pixel generation, jointly optimizing both stages with RL.
@@
RL has also been extended beyond vision–language models to a diverse range of modalities, including audio. Within the audio–language domain, we categorize RL applications into two broad classes. (1) Reasoning enhancement for large audio–language models: RL is leveraged to guide models in producing structured, step-by-step reasoning chains for tasks such as audio question answering and logical inference (Wen et al., 2025; Diao et al., 2025; Li et al., 2025d; d; Wen et al., 2025). (2) Fine-grained component optimization in speech synthesis (TTS): RL is employed to directly refine system components—for example, improving duration predictors—using perceptual quality metrics such as speaker similarity and word error rate as reward signals, thereby yielding more natural and intelligible speech (Li et al., 2025m). Some other works such as EchoInk-R1 (Xing et al., 2025) further enrich visual reasoning by integrating audio–visual synchrony under GRPO optimization.